OpenStudy (anonymous):

How do you know if a quadratic equation will have one, two, or no solutions? How do you find a quadratic equation if you are only given the solution? Is it possible to have different quadratic equations with the same solution? Explain. Provide your classmate’s with one or two solutions with which they must create a quadratic equation.

OpenStudy (saifoo.khan):

hello pixie. sorry i left yesterday bcoz my internet wasn't working and I am sorry tht u missed ur appendix

OpenStudy (anonymous):

tis ok :) i will be loosing my internet this evening, in about 2 hours, so trying to get done what i can

OpenStudy (saifoo.khan):

alright, cool.

OpenStudy (saifoo.khan):

so where are u stuck?

OpenStudy (anonymous):

the whole, my mindset today is not where it should be, and i just can't comprehend any of this today :(

OpenStudy (anonymous):

i really just need an example that i can use for my classmates to solve, i understand how to do them, but not sure about the example it asks to come up with

jhonyy9 (jhonyy9):

1. when D what is = b2 -4ac is > 0 sign that the respective quadratic equation will have two different real solutions - when D =0 the respective quadratic equation will have two solutions what will be equales - when D<0 this equation will have just two complex solutions in the form a+bi

jhonyy9 (jhonyy9):

when we have given the real solutions x_1 and x_2 yes we can recreat this quadratic equation because we know the formula x_1 +x_2 = -b/a and (x_1)*(x_2)=c/a we know that the original form og one quadratic equation is ax2 +bx +c =0 - from those formulas with name ,,la Viéte"s formulas we can calcule this a,b and c and hence we can rewrite this quadratic equation with solutions what are given

jhonyy9 (jhonyy9):

- and yes is possibille that two different quadratic equations have the same solutions